OpenStudy (anonymous):

Acceleration of a particle is increasing linearly with time. It’s acceleration is 10m/s2 at t=0 and 20 m/s2 at t=5s. Also, it’s speed is 5 m/s at t=0 and it’s position is x= -100m at t=0.Find at time t it’s acceleration, velocity and position.

OpenStudy (anonymous):

would this involve solving differential equations?

OpenStudy (anonymous):

yup..i think so

OpenStudy (anonymous):

then we got\[da /dt = 4m/s^3 \]

OpenStudy (anonymous):

\[so ......d^3x/dt^3 = 4.. \] solve this and you ll get the equation for displacement

OpenStudy (anonymous):

what's the ans u got ??

OpenStudy (anonymous):

can't you solve that D.E ?

OpenStudy (shubhamsrg):

We are given a= kt + c a=10 at t=0 and a=20 at t=5 using this, we can find out k and c. now a = dv/dt => dv = ktdt + cdt integrate both sides you'll get something like v = kt^2 /2 + ct + m v=5 at t=0, so you can find out m now v=ds/dt same way find out s=f(t). Hope this helps ?

OpenStudy (anonymous):

thanks :D it really did help :)

OpenStudy (anonymous):

:P..

OpenStudy (vincent-lyon.fr):

By the way, "its" is not "it's" !

OpenStudy (anonymous):

oh great ..thanks for correcting it...